### v2

# Praxis-VLM: Vision-Grounded Decision Making via Text-Driven Reinforcement Learning

Abstract
Vision Language Models exhibited immense potential for embodied AI, yet they often lack the sophisticated situational reasoning required for complex decision-making. This paper shows that VLMs can achieve surprisingly strong decision-making performance when visual scenes are represented merely as text-only descriptions, suggesting foundational reasoning can be effectively learned from language. Motivated by this insight, we propose Praxis-VLM, a reasoning VLM for vision-grounded decision-making. Praxis-VLM employs the GRPO algorithm on textual scenarios to instill robust reasoning capabilities, where models learn to evaluate actions and their consequences. These reasoning skills, acquired purely from text, successfully transfer to multimodal inference with visual inputs, significantly reducing reliance on scarce paired image-text training data. Experiments across diverse decision-making benchmarks demonstrate that Praxis-VLM substantially outperforms standard supervised fine-tuning, exhibiting superior performance and generalizability. Further analysis confirms that our models engage in explicit and effective reasoning, underpinning their enhanced performance and adaptability.
Figure 1: Illustrative examples of Praxis-VLM’s decision-making process. Employing text-driven training, Praxis-VLM performs sophisticated reasoning by analyzing visual situations, posing relevant questions, and generating reasoned textual responses to support multimodal decision-making.
1 Introduction
“Language is the dress of thought.” — Samuel Johnson
Developing intelligent agents capable of complex real-world interaction necessitates robust, vision-grounded situational decision-making [1, 2, 3]. Vision Language Models (VLMs) exhibited immense promise for this purpose, offering a foundation for agents that can perceive and understand visual environments [4]. However, current VLMs often lack the explicit reasoning capabilities needed to parse nuanced visual scenarios and make optimal decisions [5, 6, 7]. This limitation hinders their deployment in crucial applications, like robotics [1, 2, 3] and interactive assistants [8, 9], where the capacity to “think before decide,” much like humans do, is paramount for safe and effective operation.
Meanwhile, advancements in large language models (LLMs), such as DeepSeek-R1 [10] and OpenAI o1 [11], highlight the potential of multi-step reasoning for tackling complicated tasks. Recent efforts have attempted to enhance VLMs with sophisticated reasoning capability from text-based models [12, 13, 14]. These approaches typically utilize reasoning-oriented LLMs to generate long-form reasoning chains, which are then used to supervise VLM training. However, they rely heavily on large-scale, high-quality vision data paired with textual ground-truth answers, which are notoriously expensive and laborious to curate across diverse real-world scenarios [15, 16, 17]. The challenge of obtaining such paired image-text training data becomes even more pronounced in real-world embodied decision-making contexts with diverse situations.
This data acquisition challenge consequently motivates us to investigate the fundamental nature of decision-making abilities and their reliance on direct multimodal training. An essential question then arises: Is the core of decision-making ability exclusively tied to direct multimodal experience? If not, there may be more cost-effective pathways to improve the multimodal decision-making ability of VLMs. To address this, we conduct a preliminary analysis (§ 2) and find a surprisingly effective yet underexplored alternative: when vision-grounded situations are represented by textual descriptions, even standard VLMs could achieve comparative or even improved performance on complex multimodal decision-making benchmarks like VIVA [18] and PCA-Bench [19]. This observation sparked our central hypothesis: fundamental decision-making knowledge and reasoning capability can be learned primarily from language-based representations and then effectively transferred to visually-situated contexts during inference. This notion resonates with the mental model theory [20], which posits that humans construct internal, often language-based, representations of situations to reason, predict outcomes, and guide their decisions, later applying these models to sensory experiences.
Motivated by this insight, we conduct a systematic investigation into this hypothesis and propose Praxis-VLM, a reasoning VLM that learns high-level decision-making principles from language and applies this “praxis” within vision-grounded environments. Specifically, we begin by constructing a text-based training corpus where visual situations are articulated through descriptive language, mitigating the need for image-text paired data. Then, to foster robust and transferable reasoning—the ability to “think before decide”—we employ a Reinforcement Learning (RL) approach. Specifically, we employ GRPO algorithm [21] with multi-stage training to encourage the model to generate explicit reasoning chains before reaching a decision. To facilitate efficient learning, we introduce a novel adaptive R1 reward that targets different skills at each training stage. This learned capacity for deliberate reasoning then transfers effectively when Praxis-VLM encounters and processes actual visual inputs during multimodal inference. Illustrative examples are shown in Figure 1.
To evaluate Praxis-VLM, we adopt challenging decision-making benchmarks spanning diverse tasks: VIVA for human-centered situations, PCA-Bench for embodied agent tasks, and EgoNormia [22] for first-person video understanding. The results show that Praxis-VLM outperforms both the vanilla VLMs and SFT baselines. More importantly, it exhibits remarkable generalizability, suggesting that the reasoning skills acquired from the text are indeed fundamental and transferable. Moreover, in-depth analysis reveals that Praxis-VLM considers multiple meaningful dimensions of decision-making, such as situation analysis, consequence evaluation, safety considerations, and norm adherence (§ 5.4). This underpins both the improved decision quality and the potential for cross-domain transfer. Finally, the analysis of common reasoning errors further provides valuable insight for future research.
In summary, our work makes three-fold contributions: (1) We show the potential of leveraging language as a primary medium for instilling transferable reasoning skills in VLMs for situational decision-making; (2) Building on this, we propose a text-based reinforcement learning paradigm and introduce Praxis-VLM, a novel model that learns high-level decision-making principles from language and grounds them in concrete, multimodal scenarios via an adaptive R1 reward; (3) Through extensive experiments and analyses across three diverse benchmarks, we demonstrate Praxis-VLM’s superior reasoning and generalization capabilities for embodied decision-making, charting a cost-efficient path for scalable VLM reasoning.
2 Preliminary Analysis
The primary goal of this research is to enhance the vision-grounded situational decision-making capabilities of VLMs, enabling them to effectively reason about visually perceived situations and take appropriate actions. While recent advancements show promise in equipping models with thinking processes to tackle complex reasoning tasks [23, 24], a significant bottleneck remains: the scarcity of large-scale, annotated datasets that pair visual inputs with optimal actions and reasoning steps.
To investigate alternative pathways for developing decision-making skills , we preliminarily analyze the performance of VLMs under different input conditions. We evaluated Qwen2.5-VL [25], on two vision-grounded decision-making benchmarks: VIVA [18] and PCA-Bench [19]. They require the model to choose the best action from multiple options based on an image depicting a specific situation. We compare two settings: (1) using the original image as input situation, and (2) using a textual description of the situation, either a caption generated by GPT-4o or taken from the dataset’s annotations, in place of the image.
Figure 2: Model accuracy on VIVA [18] and PCA-Bench [19]. Image Situation uses the original image as input, and Text Situation employs the caption (text) instead.
The results, presented in Figure 2, reveal a compelling insight: the text situation setting could demonstrate performance comparable to, or even slightly better than, the VLM operating directly on the image-text paired input. This observation suggests that the fundamental reasoning and decision logic required for navigating these human-centric and embodied-agent scenarios can be substantially captured and learned from symbolic, textual representations alone.
Such findings resonate with human cognitive development, where abstract knowledge, reasoning skills, and decision-making strategies are often acquired through language, detached from an immediate sensory perception of the real situation. Based on this insight, we hypothesize that VLMs can similarly benefit from acquiring reasoning capabilities primarily through text-based learning. Therefore, our core methodological premise is to cultivate sophisticated reasoning and decision-making policies using rich, text-only situational descriptions paired with desired outcomes. The ultimate aim is to transfer these textually-learned reasoning skills effectively to multimodal inference for the model to ground its decisions in the actual visual information. It enables the model to leverage the text data for reasoning development while retaining the VLM’s ability to perceive and act in vision-grounded contexts.
3 Method: Learning to Reason and Decide from Text
Figure 3: Overview of Praxis-VLM: Learning transferable reasoning from text-only data for multimodal decision-making. The framework involves creating synthetic text-based training data, training the VLM with GRPO with adaptive rewards on this data to develop reasoning, and applying the learned reasoning to make decisions in vision-grounded situations during inference.
Our primary goal is to enhance the reasoning and decision-making capabilities of VLMs in vision-grounded scenarios. Recognizing the challenges in acquiring large-scale image-text pairs for training and inspired by §$\S$ 2, we propose a novel paradigm to learn decision-making skills from text-only data and transfer these skills to multimodal inference. Here, we employ Reinforcement Learning to foster the model’s ability to generate explicit reasoning processes for complex decision-making. As illustrated in Figure 3, our framework involves three key phases: (1) Creating a text-based decision-making dataset, (2) Optimizing the VLM’s reasoning and decision-making using GRPO [21], and (3) Deploying the enhanced VLM for decision-making with actual visual input during inference.
3.1 Problem Formulation and Model Setup
We start the methodology description with the problem formulation for vision-grounded decision-making: an agent (VLM) receives a visual situation $x^{S}$ (e.g., an image or video frame) and a textual question $x^{Q}$ about action selection. The objective is to learn a policy $\pi(y|x^{S},x^{Q})$ that generates a response $y$ maximizing task success or alignment with desired criteria (e.g., human preferences, safety constraints). We initialize the policy $\pi$ using Qwen2.5-VL [25] for its strong capabilities in multimodal understanding and instruction following, which enables a solid foundation. The VLM architecture $\mathcal{M}$ naturally supports the joint processing of visual and textual inputs: $\hat{y}=\mathcal{M}_{\theta}(x^{S},x^{Q})$.
Inspired by §$\S$ 2, importantly, our training strategy focuses on enhancing reasoning capabilities primarily through text. Therefore, during the training phase, visual inputs $x^{S}$ are replaced by their textual descriptions $x^{S}_{text}$, and only the language model components of the VLM are updated. This leverages the scalability of text data for knowledge acquisition. Yet, during the inference phase, the entire trained VLM architecture, including the vision encoder, is used to process the image-text input pair $(x^{S},x^{Q})$, allowing the textually-learned reasoning skills to be applied to visual situations.
3.2 Text-Based Decision-Making Data Construction
To gather reasoning skills, a cornerstone of our methodology is the creation of a high-quality, text-only dataset specifically designed to teach complex decision-making reasoning. This dataset serves as the primary training ground for our model learning. Its design aims to be: (1) challenging enough to necessitate multi-step reasoning for optimal decision-making, and (2) structured to allow evaluation via straightforward rule-based metrics, mitigating the need for complex reward modeling and reducing the risk of reward hacking [26]. We hence formulate the task as question answering based on a textual scenario: $(x^{S}_{text},x^{Q},y)$, where $x^{S}_{text}$ is the textual description that replaces the visual situation, $x^{Q}$ is a multiple-choice question of decision making relevant to the situation, and $y$ is the answer.
Leveraging recent advances in LLM-based data synthesis [17, 27], we employ GPT-4o [28] for generation. Our process begins with 10 manually crafted seed questions serving as in-context examples. To further maximize the diversity of the situations, we implement a batch generation approach where GPT-4o produces 10 samples per time, followed by a deduplication step. This strategy allows an effective generation of varied scenarios and questions, yielding a dataset comprising 10K training samples and 1K validation samples. More details of data creation are in Appendix A.3.
3.3 Reasoning Policy Optimization via GRPO
To cultivate robust reasoning abilities that go beyond the behavioral cloning limitations of supervised fine-tuning (SFT) [29], we employ Reinforcement Learning (RL) to fine-tune the VLM’s policy. Specifically, we utilize Group Relative Policy Optimization (GRPO) [21], an RL algorithm well-suited for optimizing decision-making policies based on sampled trajectories to enhance reasoning.
Concretely, given an old policy $\pi_{old}$ and a reference policy $\pi_{ref}$, GRPO optimizes the current policy $\pi_{\theta}$ by sampling $G$ response trajectories $\mathcal{O}_{i}=\{o_{i}\}_{i=1}^{G}$ for each query $x$. The objective function is:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{D},\{\mathcal{O}_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot|x)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\min\Bigg{(}\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{old}}(o_{i}|x)}\hat{A}_{i},\\ \text{clip}\Big{(}\frac{\pi_{\theta}(o_{i}|x)}{\pi_{\text{old}}(o_{i}|x)},1-\epsilon,1+\epsilon\Big{)}\hat{A}_{i}\Bigg{)}-\beta D_{\text{KL}}[\pi_{\theta}(\cdot|x)\parallel\pi_{\text{ref}}(\cdot|x)]\Bigg{]}$ (1)
where $\epsilon>0$ is a policy ratio clipping hyperparameter, $\beta>0$ balances the KL-penalty term against the advantage-weighted policy update, and $D_{KL}[\pi_{\theta}||\pi_{ref}]$ is the KL divergence between the current and reference policy. The term $\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},...,r_{G}\})}{\text{std}((\{r_{1},r_{2},...,r_{G}\})}$ represents the normalized advantage estimate of the $i$-th response at the group level. GRPO aims to improve the policy by increasing the likelihood of actions that lead to higher-than-average estimated returns within a sampled group.
3.4 Multi-Stage GRPO Training with Adaptive R1 Reward
To further encourage robust, explicit reasoning capabilities, we employ multi-stage GRPO training. It is inspired by recent findings that geometry and math data can enhance model logical reasoning ability [10, 14, 30]. Multi-stage GRPO, combined with a newly designed adaptive R1 reward tackling different aspects [31], allows us to first establish foundational logical structuring and then refine sophisticated decision-making skills, enabling models to learn different skills at different stages.
Stage 1: Cold Start Initialization for Foundational VLM Reasoning. The initial stage focuses on equipping the VLM with multi-step reasoning abilities. We employ the geometry3k dataset [32] for GRPO training, converting the task into numerical computation which can be readily evaluated using rule-based metrics. Following DeepSeek-R1, we enforce a specific output format: "<think> </think><answer> </answer>", compelling the model to externalize its reasoning process.
A key finding of our work is that the commonly adopted SFT-then-RL paradigm, where a model is first fine-tuned on (image, question, reason) triplets with SFT to learn the desired reasoning structure before RL [12, 13], can be circumvented. We find that directly training an instruction-tuned VLM (e.g., Qwen2.5-VL-Instruct) with GRPO is effective when coupled with an adaptive reward mechanism. In this initial phase, the rewards prioritize format adherence. Specifically, we leverage a combination of: (1) $R_{\text{tag}}$: Calculates if the count of each special token (<think>, </think>, <answer>, </answer>) in the output equals one, which strongly encourages the model to learn the basic output structure and narrows the search space; (2) $R_{\text{format}}$: Measures if the output strictly adheres to the exact nested format; (3) $R_{\text{accuracy}}$: Rewards the correctness of the final numerical answer.
Once the model consistently produces outputs in the desired format, $R_{\text{tag}}$ is removed, and the focus shifts more towards $R_{\text{accuracy}}$, thereby promoting better reasoning and result accuracy.
Stage 2: RL Training for Text-Based Decision Making. The model emerging from Stage 1, now possessing a better-initialized capability for multi-step reasoning and format adherence, serves as the foundation for the second training stage. This stage targets our primary goal: enhancing sophisticated decision-making skills. Here, we utilize the curated text-based decision-making dataset and train the model to mimic human-like learning processes by exploring diverse reasoning paths for various textual scenarios. The reward function in this stage primarily emphasizes the correctness of the final decision, implicitly validating the quality of the preceding reasoning. Leveraging text data in this manner allows for the efficient adaptation and refinement of reasoning skills for sophisticated decision-making, ultimately yielding a policy designed for effective transfer to multimodal inference.
For this decision-making RL training, the adaptive rewards aim to encourage both comprehensive thinking and accurate decisions. We use a combination of: (1) $R_{\text{format}}$: Continues to ensure adherence to the thinking-answering structure; (2) $R_{\text{accuracy}}$: Rewards the correctness of the answer; (3) $R_{\text{len}}$: Encourages the model to generate deliberate and longer reasoning chains. Our observations indicate that $R_{\text{len}}$ is effective in promoting a more comprehensive consideration of the situation and action candidates. Contrary to some recent work suggesting that long reasoning chains might be redundant [33, 34], our results show that encouraging longer reasoning can lead to more thorough situational analysis for more complex situations. We will provide detailed discussions on this in §$\S$ 5.2. This adaptive reward strategy across stages enables efficient training by targeting different skills, including format adherence, logical computation, and complex decision-making, sequentially.
4 Experimental Settings
4.1 Tasks and Datasets
To comprehensively assess the decision-making capabilities of our model in diverse vision-grounded scenarios, we utilize three benchmarks that encompass a wide spectrum of real-world situations.
VIVA [18]: This benchmark focuses on human-centered situations. It comprises 1,240 images depicting a variety of real-world scenarios. Models are tasked with understanding social contexts and predicting appropriate actions or responses aligning with human values based on the visual scenes.
PCA-Bench [19]: A benchmark encompassing 317 scenarios of embodied robotics, autonomous driving, and interactive games. It requires models to process multimodal observations and select an action from a predefined action space. We use the open track proportion of the benchmark.
EgoNormia [22]: A dataset with 1,743 samples centered around ego-centric video understanding, where the model needs to interpret actions and anticipate future events from an egocentric perspective.
We consider VIVA and PCA-Bench as in-domain benchmarks, as they align with the typical image-text input and decision-making formats our model is trained on. In contrast, EgoNormia serves as an out-of-domain benchmark, introducing additional challenges such as sequential and temporal reasoning over video frames and egocentric perception. These datasets offer a rigorous and diverse testbed for evaluation. We follow the original data splits and prompts provided by each benchmark.
4.2 Baselines and Implementations
Following previous work in reasoning-based VLMs [13, 35], we adopt Qwen2.5-VL as our backbone model, with both its 3B and 7B parameter variants. We compare the performance of our model (Praxis-VLM) against baselines, including original backbone (vanilla) VLMs and the SFT method. The SFT baselines include two variants: one (w/ SFT) that directly predicts the answer $y$, and another (w/ Reason SFT) that first generates a reasoning chain before producing the final answer.
Implementation Details. For both GRPO and SFT training, we finetune full model parameters. For GRPO training, we set rollout N to 5 and KL divergence coefficient to 0.01. During inference, we leverage VLLM Library [36] with greedy decoding. More details are in Appendix A.2.
5 Results and Analysis
Models VIVA [18] PCA-Bench [19] EgoNormia [22]
Qwen2.5-VL-3B 76.61 48.58 51.92
       $\hookrightarrow$ w/ SFT 77.42 46.37 35.06
       $\hookrightarrow$ w/ Reason SFT 75.81 49.53 28.34
Praxis-VLM-3B (ours) 79.03 50.79 54.27
       $\hookrightarrow$ w/ one-stage GRPO 79.52 50.79 53.13
Qwen2.5-VL-7B 80.97 46.37 46.19
       $\hookrightarrow$ w/ SFT 81.13 45.74 34.83
       $\hookrightarrow$ w/ Reason SFT 78.79 53.00 34.08
Praxis-VLM-7B (ours) 84.03 60.25 54.33
       $\hookrightarrow$ w/ one-stage GRPO 83.87 58.99 49.57
Table 1: Main results measured by accuracy (%). w/ SFT denotes the SFT baseline to directly predict the answer, while w/ Reason SFT first generates a reasoning chain before producing the answer. w/ one-stage GRPO is our model variant without math cold start initialization (Stage 2 only).
5.1 Main Results
The main results are in Table 1, showing several key advantages of the proposed Praxis-VLM. First, our text-based GRPO training strategy effectively endows VLMs with robust decision-making skills that successfully transfer to multimodal scenarios. Across both 3B and 7B model scales, Praxis-VLM consistently outperforms Qwen2.5-VL-Instruct as well as SFT approaches on all benchmarks. This primary observation underscores the core efficacy of our approach: it successfully imbues the model with decision-making capabilities learned from textual scenarios, which are then effectively transferred and applied during multimodal inference in varied visual environments.
Second, Praxis-VLM exhibits superior generalization ability compared to SFT-based approaches. This advantage is particularly salient in the out-of-domain EgoNormia dataset, which features sequential video inputs distinct from our training regime. Praxis-VLM maintains strong performance here, a stark contrast to both SFT baselines, which struggle significantly when faced with such domain shifts. Such a disparity suggests that while SFT-based methods primarily learn to imitate the patterns seen during training with behavioral cloning, our GRPO-trained model internalizes more fundamental and broadly applicable decision-making skills. In contrast, while Reason SFT learns to replicate the reasoning patterns seen during training, it appears to overfit these specific patterns and struggles to adapt its reasoning when faced with domain shifts. By optimizing the policy based on task outcomes and allowing exploration beyond static dataset examples, Praxis-VLM learns to analyze situations, evaluate potential actions, and understand consequences in its own generated reasoning paths, cultivating a generalizable decision-making competency across diverse situations.
Third, the cold-start initialization in our multi-stage framework further enhances the model’s generalization capabilities, particularly for novel and complex tasks. Comparing the full, two-stage Praxis-VLM with its one-stage variant (without the initial math cold start) reveals a distinct improvement in generalization: while both variants achieve comparable performance on in-domain tasks (VIVA, PCA-Bench), the two-stage Praxis-VLM consistently exhibits superior accuracy on the EgoNormia benchmark. This demonstrates that the math cold-start successfully bolsters the model’s foundational logical reasoning architecture, thereby enhancing its capacity to adapt and perform effectively in novel and complex decision-making scenarios. Besides, the performance gain is more pronounced in 7B models, possibly because of larger models’ better logical reasoning potential.
In summary, the results demonstrate that Praxis-VLM , enhanced with text-based GRPO training, achieves substantial improvements in vision-grounded decision-making. It effectively leverages textual guidance to learn generalizable decision-making capabilities, enabling robust performance across diverse complex multimodal scenarios. Compared to SFT approaches and the original base models, Praxis-VLM exhibits markedly stronger generalization and adaptability.
5.2 Impact of Reasoning Length on Model Performance
Figure 4: Accuracy versus reasoning length on VIVA and EgoNormia. Samples are grouped into 5 quintile bins based on the reasoning length percentile generated by Praxis-VLM (Len1: shortest 20%, Len5: longest 20%).
Previous results have shown explicit reasoning helpful to our task. For a further analysis, we measure the length of the generated reasoning chain produced for each sample in VIVA and EgoNormia.  [^1] We do not include PCA-bench due to the relatively small sample sizes. Based on these lengths, we divide the samples into five equal bins by length percentiles. We then calculate the accuracy within each bin for both Praxis-VLM and Qwen2.5-VL. The results are presented in Figure 4.
We can observe a general trend of decreasing accuracy for Praxis-VLM as the reasoning length increases for both datasets. However, crucially, the accuracy of the baseline Qwen-VL, without reasoning, also shows a similar downward trend. This strongly suggests that the decreasing accuracy is correlated with sample difficulty; Praxis-VLM tends to generate longer, more detailed reasoning for instances that are inherently more challenging.
Moreover, within nearly every length bin, Praxis-VLM consistently outperforms its corresponding Qwen2.5-VL baseline. This holds true for both the 3B and 7B models. This finding further reinforces the effectiveness of the explicit reasoning process learned by Praxis-VLM, demonstrating its robust benefit across varying levels of sample complexity.
Finally, we observe a noticeable performance drop for Praxis-VLM specifically in the longest reasoning bin (Len5) on the VIVA dataset than Qwen-VL. We then manually examine these cases and find two potential contributing factors. First, some generated outputs exceed the maximum sequence length configured during inference (i.e., 1,024 tokens), causing the generation to be truncated before the final answer tag could be produced. Second, extremely long reasoning chains might sometimes cause “overthinking,” where the extended reasoning process potentially introduces noise, negatively impacting the final decision accuracy. This point may warrant further investigation in future work.
5.3 Diverse Reason Sampling for Enhanced Decision Making
To further evaluate the robustness of Praxis-VLM’s decision-making, we generate 8 diverse samples per instance with decoding temperature as 0.2 and measure accuracy via Majority Vote ("Major.", most frequent answer) and Pass@1 (at least one correct answer). The results are shown in Table 2.
Compared to greedy decoding (“Orig.”), we observe that for reasoning-enhanced models like Praxis-VLM and Reason SFT, the majority vote yields improved accuracy. More significantly, the Pass@1 scores demonstrate substantial improvement scores. This indicates that while the single most probable reasoning path might not always lead to the correct answer, the correct solution is often reachable and present within a small set of diverse reasoning trajectories explored by these models.
Model Name VIVA PCA-Bench EgoNormia
Orig. Major. Pass@1 Orig. Major. Pass@1 Orig. Major. Pass@1
Qwen2.5-VL-7B 80.97 80.73 80.81 46.37 48.27 56.47 46.19 46.36 54.50
   w/ SFT 81.13 81.21 83.55 45.74 46.37 50.16 34.83 34.60 40.79
   w/ Reason SFT 78.79 80.64 89.03 53.00 58.36 82.33 34.08 35.69 66.04
Praxis-VLM-7B 83.87 84.36 89.27 58.99 61.83 77.92 49.57 55.08 72.23
Table 2: Performance of diverse sampling. Orig.: Greedy decoding accuracy. Major.: Majority vote accuracy with 8 distinct samples. Pass@1: Accuracy with at least one correct answer from 8 samples. For Praxis-VLM, we use the one-stage variant without math cold start for a fair comparison with SFT.
Moreover, we can observe that despite Reason SFT’s ability to sometimes find the correct answer within its samples (high Pass@1), Praxis-VLM consistently outperforms it in the Majority Vote metric across all datasets. This suggests that while both models explore relevant reasoning paths, the central tendency of Praxis-VLM’s reasoning (as reflected by the majority vote) is more reliably accurate. We interpret this as evidence for a higher quality or more robust reasoning process learned via GRPO. Overall, the results highlight Praxis-VLM’s strength in both exploring the solution space effectively (high Pass@1) and converging on the correct answer (strong Major. and Orig.).
5.4 Exploring Praxis-VLM’s Reasoning: What Does It Consider?
Figure 5: Dominant reasoning dimensions used by Praxis-VLM in decision-making. Clusters were identified by analyzing keyphrases generated by GPT-4o from the model’s reasoning chains.
To gain deeper insights into the reasoning capabilities learned by Praxis-VLM, we analyze its generated reasons. We first prompt GPT-4o to generate keyphrases that summarize the reasoning aspects of each sample. These keyphrases are then clustered across the dataset. As shown in Figure 5, this analysis reveals four primary dimensions characterizing the model’s reasoning process.
A major aspect identified is Situational Analysis, which focuses on interpreting the scenario, assessing visual evidence, recognizing the core problem, and establishing situational awareness. This also highlights that the model can adapt the reason to multimodal inputs. Complementing this understanding is Action & Consequence Evaluation, where the model systematically compares potential actions, anticipates both positive and negative outcomes, weighs trade-offs, and often employs elimination strategies to justify its decisions. Another key dimension is Safety & Risk Management, where the model actively identifies potential hazards, considers risk mitigation strategies, and prioritizes health and safety, indicating its potential in human-centered consideration. Lastly, reasoning also incorporates Rule & Norm Adherence, which entails consideration of explicit regulations, implicit social or cultural norms, procedural correctness, and the appropriateness of involving authorities.
Taken together, the clustering results suggest that Praxis-VLM develops a comprehensive and structured approach to reason in decision-making. It systematically analyzes situations, deliberates over potential actions and their consequences, and accounts for key constraints related to safety, rules, and norms. This multifaceted capability, cultivated through text-based RL, underpins the model’s improved performance and its ability to generalize across diverse scenarios.
5.5 Understanding Failures: Error Analysis Through the Lens of Reasoning
We have shown Praxis-VLM’s superiority, and here we discuss its limitations by analyzing the reasoning chains associated with error samples. This qualitative analysis helps pinpoint common failure modes in the model’s reasoning process, even when it attempts to deliberate step-by-step. Figure 6 illustrates examples of the common error categories identified through this analysis.
First, Misinterpretation of Situation Context remains a challenge. In these cases, the model might acknowledge certain details but fail to grasp their full implication or overlook other critical contextual elements, leading to a flawed assessment of the situation (e.g., In the image of Example 1, the model fails to recognize the actual green light indicating permission to proceed straight). Second, we observed Safety & Prioritization Failures. This category includes errors where the model struggles to appropriately weigh immediate actions against long-term safety needs or fails to prioritize the most critical safety concern among multiple factors present in the scenario (e.g., in Example 2, the model addresses an immediate risk without selecting the best overall preventative measure). Third, errors frequently arise from Normative & Appropriateness Misalignment. Here, the model may make questionable assumptions about social interactions or select actions based on flawed reasoning about social norms, ethics, or expected behavior in a given context (e.g., in Example 3, the model fails to understand that driving under the influence is illegal and should be stopped).
Figure 6: Illustrative examples of common failure cases for Praxis-VLM, identified by analyzing the reasoning generated for incorrect predictions. Complete outputs are in Appendix B.
These failure modes highlight that while encouraging explicit reasoning is beneficial, challenges remain in ensuring deep and accurate contextual understanding, robust prioritization under complex constraints, and reliable alignment with nuanced social and ethical norms. Addressing these aspects within the reasoning process is a key direction for future research. More samples are in Appendix B.
6 Related Work
VLMs in Embodied Decision-Making. The quest to enable intelligent agents to make informed decisions in real-world, situated environments is central to embodied AI research [2, 22, 37, 38]. VLMs have emerged as a powerful foundation for such agents, demonstrating significant promise in applications like robotics, autonomous navigation, and interactive task planning [39, 40, 41, 42]. These models integrate visual perception with language understanding to interpret and interact with their surroundings. However, a persistent challenge lies in equipping VLMs with the capacity for multi-step reasoning required to navigate and act effectively in nuanced and dynamic situations. Our work focuses on enhancing such sophisticated reasoning abilities crucial for robust decision-making.
Reasoning in VLMs. The ability to reason is fundamental to effective decision-making. Traditional methods often rely on fine-tuning VLMs on large multimodal datasets tailored to specific tasks or reasoning styles [16, 43, 44]. More recently, approaches have emerged that encourage VLMs to generate explicit reasoning steps, often leveraging RL to optimize performance on complex tasks [13, 35, 45, 46, 47]. While these RL-based methods have shown success in improving reasoning, they typically necessitate extensive training on datasets comprised of paired image-text data. Our text-driven RL diverges from this by proposing a more data-efficient pathway to instill reasoning.
Text-Driven Enhancement of VLMs. Leveraging textual data to enhance VLMs is an area of growing interest. Some prior studies have utilized text to improve VLMs by aligning the embedding spaces of different modalities  [48, 49, 50]. However, these approaches generally do not target the cultivation of sophisticated reasoning abilities for situated decision-making. Building on our preliminary analysis, we introduce a novel method to employ text-driven RL to instill a generalizable decision-making competency. A crucial aspect of our contribution is finding that these reasoning skills, learned entirely from text, can be effectively transferred to diverse vision-grounded scenarios for decision-making.
7 Conclusion
We introduce Praxis-VLM, a reasoning-based VLM for complex vision-grounded decision-making. Motivated by our finding that foundational reasoning can be effectively learned from text-only descriptions, Praxis-VLM utilizes text-based GRPO to instill robust reasoning skills that successfully transfer to vision-grounded inference. The experiments on three benchmarks of diverse situations demonstrate that our approach outperforms the original VLMs and different SFT methods, particularly in generalization to out-of-domain tasks with general reasoning abilities. Our work offers a data-efficient pathway to more capable and generalizable VLMs by effectively transferring abstract reasoning learned from language to guide complex vision-grounded decision-making.
Discussions on Limitations and Future Work
The explorations in this work with Praxis-VLM open up several exciting avenues for future research, offering insights into how abstract reasoning learned from language can be effectively grounded in multimodal contexts. While we demonstrated significant gains using 3B and 7B parameter models, a broader investigation into the interplay between model scale and the efficacy of text-driven reasoning transfer would be beneficial. Understanding how models of different size perform could reveal valuable scaling dynamics for this learning paradigm.
Furthermore, our current approach leverages a curated corpus of text-only situational descriptions. An insightful direction for future work lies in optimizing this data aspect further. While text offers a data-efficient route to learning reasoning, exploring advanced data selection strategies could unlock even greater efficiency. This could involve identifying or generating a smaller subset of highly "effective" textual scenarios that most potently instill transferable reasoning skills.
Another promising frontier involves enhancing the synergy between the text-learned reasoning and the VLM’s foundational visual perception, which, however, is out of the scope of our work. The ultimate effectiveness of the transferred reasoning during multimodal inference hinges on how accurately the visual input is perceived and aligned with the conceptual understanding developed through text. Future research could focus on improving the VLM’s core visual grounding capabilities, perhaps through targeted pre-training or co-training strategies that explicitly link visual features to the abstract reasoning structures learned from language.
Finally, despite the effectiveness of our method, the error analysis shows several common fail pattens Praxis-VLM tends to exhibit during the reasoning process. Future work may address these issues to further enhance the model’s performance.
Ethics Statement and Broader Impacts
The evaluation of our models in this research is conducted using publicly available benchmarks, including VIVA, PCA-Bench, and EgoNormia. We adhered strictly to the original usage protocols and licensing terms of these benchmarks, utilizing them without modification and solely for model inference during the evaluation phase. For the generation of any synthetic text-based training data using LLMs like GPT-4, we employ keyword-based filtering mechanisms designed to mitigate the inclusion of potentially harmful, biased, or unsafe content.
Despite these precautions, it is important to acknowledge that our work, Praxis-VLM, builds upon pre-trained VLMs. These foundational models are typically trained on extensive datasets scraped from the internet, which may inadvertently contain and reflect existing societal biases or problematic content. While our method focuses on learning reasoning skills, the potential for the model to inherit or amplify such underlying issues from its base architecture remains. We therefore strongly advise users and developers to conduct thorough ethical reviews, bias assessments, and impact analyses before deploying systems based on this research in real-world applications, particularly in sensitive or high-stakes domains.
Appendix A Experiment Details
A.1 Data Statistics
Task Number
VIVA 1,240
PCA-Bench 317
EgoNormia 1,743
Table 3: Data statistics for each evaluation benchmark. "Number" refers to the count of test instances used.
Our experiments utilize three established benchmarks for embodied decision-making, providing a comprehensive evaluation of our model’s capabilities across diverse scenarios. Key statistics for these benchmarks are presented in Table 3. The benchmarks are: (1) VIVA [18], which is focused on human-centered decision-making, presenting various real-world social situations where the model must predict appropriate human actions; (2) PCA-Bench [19] encompasses scenarios from embodied robotics, autonomous driving, and interactive games. For our experiments, we use the open track test set provided by the benchmark; (3) EgoNormia [22], which centers on normative decision-making from an ego-centric video perspective, requiring models to interpret actions and anticipate events involving tool use or object manipulation.
Across all benchmarks, we utilize the task of action selection to measure the model decision-making ability. This is formalized as a multiple-choice question answering task, where the model is presented with a visual situation and must choose the most appropriate action from several candidates. For EgoNormia, which uses video input, we adhere to the method described in the original work: video frames are sampled at a rate of one frame per second and are then concatenated from left to right (LTR) to form a single composite image input for the model.
A.2 Implementation Details
For both GRPO and SFT training, we finetune full model parameters with BFloat16. For GRPO implementation, we use the EasyRL Library  [^2] https://github.com/hiyouga/EasyR1/tree/main. We adopt the default hyper-parameters, and set rollout N to 5 and KL divergence coefficient as 0.01. The learning rate is set as 1e-6.
For SFT implementation, we employ the HuggingFace TRL  [^3] https://huggingface.co/docs/trl/en/index. We set the number of training epochs as 3 and learning rate as 2e-5. For Reason-SFT baseline, as there is no reasoning chains available, we first utilize GPT-4 to generate a plausible reasoning chain for each textual training sample. We then fine-tune the base VLM using SFT on these augmented (Situation, Question, Reason, Answer) samples, specifically training the model to first generate the reasoning chain and then the final answer, mimicking the desired output format.
All models are trained on four NIVIDA A100 and H100 GPUs. For Praxis-VLM, we adopt the following system prompt:
System Prompt You are a helpful AI Assistant, designed to provided well-reasoned and detailed responses. You FIRST think about the reasoning process as an internal monologue and then provide the user with the answer. The reasoning process MUST BE enclosed within <think> and </think> tags, and the final answer MUST BE enclosed within <answer> and </answer> tags.
Druing inference, we leverage VLLM Library [36] with greedy decoding. The model performance is evaluated with accuracy. To parse the model output and match it to the original options (e.g., A/B/C/D/E, etc), we first design a list of rules for matching; if it cannot be matched, we prompt GPT4-o to match the model answer with the options. The prompts for each benchmark in inference are shown below:
Inference Prompt for VIVA You are given a situation and a question. Based on the situation provided, select the most appropriate option to answer the question:
## Situation: As shown in the given image. ## Question: _question_
Now answer the question. Just output the choice:
Inference Prompt for PCA-Bench Please answer the question below based on the images.
## Question: _question_
Now answer the question by selecting the correct option.
Inference Prompt for EgoNormia The given images from a first-person perspective video depict a person in a given situation. Please answer the question below based on the images.
## Question: Given the below list of behaviors, choose the single most normatively relevant or appropriate action to perform next. You shouldn’t use the info in options to learn about the context, but rather to make a decision based on the normative appropriateness of the behavior. You shouldn’t eliminate any options only based on the presence of elements in the context; you should focus on normative appropriateness. _question_
Now answer the question by selecting the correct option.
Reward. For GRPO training, we adopt rule-based rewards, eliminating the needs for parameterized reward models. For geometry3k data training, we convert the task into numerical computation and use the Math-Verify Library  [^4] https://github.com/huggingface/Math-Verify to calculate the binary reward: 1 for correct and 0 for incorrect. For text-based decision making, which is formulated as multiple-choice question, we parse the model answer using rules and match it with correct answer using a similar binary reward. For length reward ($R_{\text{len}}$), we first count the number of words in a multi-step reason, and then calculate the score as this word count divided by a scaling factor of 250. To prevent disproportionately long outputs from dominating, $R_{\text{len}}$ is capped at a maximum value of 1.0, which is achieved if the word count reaches or exceeds 250 words. For geometry3k training, the overall reward is: $R=R_{\text{accuracy}}+R_{\text{format}}+0.5\cdot R_{\text{tag}}$. For text-based decision making training, we remove the tag reward and include the length reward, with the overall score as: $R=R_{\text{accuracy}}+0.8\cdot R_{\text{format}}+0.5\cdot R_{\text{len}}$.
Figure 7: Training dynamics of Praxis-VLM-7B (one-stage GRPO).
Training Dynamics. Figure 7 illustrates key aspects of the training dynamics during the GRPO phase, including the evolution of various reward components and the mean output length. The plots demonstrate that the model initially prioritizes and rapidly masters the desired output format. Rewards associated with structural correctness, such as ‘tag count‘ and ‘format‘ (reflecting proper use and placement of reasoning tags like ‘<think>‘ and ‘<answer>‘), are observed to quickly saturate at or near their maximum values early in the training process. Once format adherence is achieved, the ‘accuracy‘, which measures the correctness of the final decision, exhibits a consistent upward trend. This signifies that the model then progressively refines its reasoning capabilities to improve decision-making accuracy. Concurrently, the output length typically shows an initial increase as the model learns to generate more detailed reasoning, followed by stabilization. These dynamics effectively demonstrate how the adaptive reward strategy guides the model: first to learn the task structure, and subsequently to enhance the substantive quality and accuracy of its reasoning and outputs.
A.3 Details on Text-Based Data Creation
Figure 8: An example of a synthetic data sample used for our text-based training. It comprises a textual situation description, a multiple-choice decision-making question, and the target answer. The reasoning chain, also generated by GPT-4o, is used for Reason SFT training, and not utilized during our GRPO training phase.
For the construction of our text-based decision-making dataset, we utilize GPT-4o 555gpt-4o-2024-11-20. A primary objective is to generate training data that is both sufficiently challenging to necessitate multi-step reasoning and diverse in its situational contexts to enhance model generalizability. To promote this situational diversity, we implement a batch generation strategy: GPT-4o is prompted to produce 10 samples per request. These generated batches subsequently underwent a deduplication process to ensure a varied collection of scenarios and questions. This method proves effective in creating the desired dataset characteristics. The prompt used for this data generation is shown below:
Prompt for Training Data Generation Now your task is to create complex decision- making questions in human-centered situations. Each question contains a situation description, a multiple-choice question, and an answer. You can consider the following approaches to enhance the complexity:
- Add more context to the problem, such as tools, back-ground information, or character details, making the constraints more specific;
- Make the options challenging;
- Consider different ways the question is asked, incorporating reverse reasoning, dialectical reasoning, critical thinking, etc.
The question doesn’t necessarily have to ask which action is correct but could focus on other aspects related to decision-making.
There are no specific format or wording requirements for the questions, but they should be in the form of multiple-choice questions. You should make the situation diverse. You should also include a rationale to explain the answer.
## Examples:
_example_
Now generate 10 candidate question with answer. Your output should be presented as a JSON list
We provide a synthetic sample for text-based training in Figure 8.
Appendix B Sample Outputs
We present error outputs of Praxis-VLM-7B with full reason paths in Figure 9. We further include model outputs on samples from VIVA, PCA-Bench, and EgoNormia in Figure 10, Figure 11, and Figure 12.
Figure 9: Error sample outputs of Praxis-VLM-7B with full model reasons.
Figure 10: Output of Praxis-VLM-7B with full model reasons on a sample from VIVA.
Figure 11: Output of Praxis-VLM-7B with full model reasons on a sample from PCA-Bench.
Figure 12: Output of Praxis-VLM-7B with full model reasons on a sample from EgoNormia.